OpenStudy (anonymous):

x-y=2 need points to graph

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

set y equal to x

OpenStudy (anonymous):

i get confused

OpenStudy (anonymous):

first you need to the equation in the formate y=mx+c

OpenStudy (anonymous):

so y=-2+x

OpenStudy (anonymous):

ok, im following

OpenStudy (anonymous):

or in y=mx+c format its y=x-2

OpenStudy (anonymous):

you can see its a linear equation !!

OpenStudy (anonymous):

can I put 1 in place of y?

OpenStudy (anonymous):

do you mean so you can start plotting a graph?? y=1 ?? is that what you mean??

OpenStudy (anonymous):

yes, I need points to graph from x-y=2

OpenStudy (anonymous):

ok. NO star x|y table. puting x vales and filling y vales using the equation y=x-2

OpenStudy (anonymous):

so for example ... let x be -2 what will y be???

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

??

OpenStudy (anonymous):